You are given the system of linear equations

"2x+ky=5, \quad x+3y=7,"

where k is a constant.


The system above has no solution when k= _________________




1
Expert's answer
2021-08-31T14:09:08-0400

The given system of linear equations is:

"2x+ky=5\\x+3y=7"

The given system is of the type:

"a_1x+b_1y+c_1=0,\space a_2x+b_2y+c_2=0"

which has no solution if:

"\frac{a_1}{a_2}=\frac{b_1}{b_2}≠\frac{c_1}{c_2}"


Therefore, the given system can be written as:

"2x+ky−5=0\\x+3y−7=0"

which has no solution if:

"\frac{2}{1}=\frac{k}{3}≠\frac{−5}{−7}\\⇒\frac{2}{1}=\frac{k}{3}⇒\\k=6"

so, the given system has no solution for k=6.